Rectangle has perimeter of 490mm, lengths sides are the ratio 3:4, the dimensions of the rectangle?

Answer by checkley79(3341)

LET THE LENGTH BE 3X & THE WIDTH BE 4X
P=2L+2W
490=2*3X+2*4X
490=6X+8X
490=14X
X=490/14
X=35
3*35=105 FOR THE LENGTH.
4*35=140 FOR THE WIDTH.
PROOF:
490=2*105+2*140
490=210+280
490=490